# The Curious Case of Arbitrariness in Machine Learning

### v1

@@ Abstract.
Algorithmic modelling relies on limited information in data to extrapolate outcomes for unseen scenarios, often embedding an element of arbitrariness in its decisions. A perspective on this arbitrariness that has recently gained interest is multiplicity—the study of arbitrariness across a set of “good models”, i.e., those likely to be deployed in practice. In this work, we systemize the literature on multiplicity by: (a) formalizing the terminology around model design choices and their contribution to arbitrariness, (b) expanding the definition of multiplicity to incorporate underrepresented forms beyond just predictions and explanations, (c) clarifying the distinction between multiplicity and other traditional lenses of arbitrariness, i.e., uncertainty and variance, and (d) distilling the benefits and potential risks of multiplicity into overarching trends, situating it within the broader landscape of responsible AI. We conclude by identifying open research questions and highlighting emerging trends in this young but rapidly growing area of research.
isbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
@@
Machine learning attempts to approximate the complexities of the world, inevitably simplifying or generalizing aspects of reality and failing to fully capture its nuances (Singh, 1993; Box, 1976; Barocas et al., 2023; Hooker, 2021; Buolamwini and Gebru, 2018; Birhane, 2022). It is thus inherently susceptible to arbitrariness, as it attempts to extrapolate outcomes based on limited information. Whether due to imperfect data (Buolamwini and Gebru, 2018; Wang et al., 2024; Geirhos et al., 2020), flawed modelling assumptions (Hooker, 2021; Breiman, 2001; Jacobs and Wallach, 2021), or the unpredictability of certain tasks (Wang et al., 2024; Rescher, 1997; Dressel and Farid, 2018), this arbitrariness is an unavoidable byproduct of any data-driven learning, including machine learning. Hence, recognizing and understanding this arbitrariness is crucial for developing responsible learning models.
@@
The study of arbitrariness is not new; it has long been a subject of interest in uncertainty literature, with roots going back centuries in statistics and decision theory (Savage, 1972; de Moivre, 1718; Bayes, 1763; Box and Tiao, 1973; Breiman, 2001; Gal, 2016; Kendall and Gal, 2017). Recently, however, a new paradigm called multiplicity has emerged (Breiman, 2001; Marx et al., 2020; Black et al., 2022). First articulated by Breiman (2001), multiplicity has gained popularity due to its unique focus only on the arbitrariness present within a set of “good models”, i.e., models that pass certain selection criteria and thus are likely to be deployed, commonly known as the Rashomon set. Moreover, multiplicity takes an intriguing perspective on the question of arbitrariness in model decisions, by instead examining arbitrariness in model selection. Through choices made during development and model selection, multiplicity offers an operational lens to the issue of arbitrariness and lays the groundwork for practical solutions in real-world applications.
Several existing works in the literature have provided broad overviews of the field of multiplicity. Black et al. (2022) holds a special place in modern multiplicity research, offering a comprehensive discussion of “opportunities”, “concerns”, and potential “solutions” of multiplicity. Similarly, Rudin et al. (2024) presents an excellent discussion on the benefits of multiplicity, with a focus on its role in identifying simpler and more interpretable models. At this point, it would be remiss not to acknowledge the PhD dissertations of Semenova (2024); Zhong (2024); Cooper (2024); Hsu (2023); Black (2022); Watson-Daniels (2024); Hasan (2022), contributing valuable perspectives on the role of multiplicity in machine learning. Despite these notable contributions, the field still lacks a systematic review of the literature–clearly needed given its rapid growth in recent years (Figure 1). To address this gap, we present the first systematic literature review of multiplicity in machine learning, consolidating existing discussions and identifying overarching trends.
To ensure comprehensive coverage of the literature, we search across various online repositories (DBLP & ACM Digital Library) using multiple search terms (‘rashomon’, ‘model multiplicity’, ‘set of good models’), followed by rigorous manual filtering. We were eventually left with 80 papers that deeply engaged with multiplicity as a central theme in their contributions. Each paper was then manually tagged with all applicable tags, and some relevant statistics are presented in Figure 1. The growing interest in the field is evident, with literature on a wide range of problems related to multiplicity. Precise details about the review process are delegated to the Appendix (§A).
Building on the insights from our review, we make the following contributions to the multiplicity literature. First, we revisit the Rashomon effect, emphasizing the role of developer choices and formalizing the terminology to trace the emergence of various subdomains within multiplicity (§2). Expanding on this discussion, we extend the existing definitions of Rashomon sets and multiplicity to include underrepresented works, making both metrics of model selection and multiplicity more intentional (§3). Next, we distinguish multiplicity from related concepts such as uncertainty and variance; and provide both mathematically grounded differences as well as practical guidance on when to adopt each perspective (§4). Finally, we explore two overarching trends in the multiplicity literature: its role in exploring diverse interpretations during model selection (§5), and its broader implications within responsible AI (§6). We conclude by identifying open research questions and future directions to encourage continued exploration in this domain.
¡2020 2020 2021 2022 2023 2024 001010101020202020303030304040404050505050Year g Number of Papers
BM&E HM REML RSE RAI App Sur PM EM OM 001010101020202020303030304040404050505050Tags
Exploring Alternative InterpretationsResponsible AIMultiplicityRashomon EffectOthers
@@
Figure 1. Statistics from the systematic review, showing the number of papers over the years and their categorization. Each paper can have multiple tags, marking all categories of contributions made by the paper. Details of the tags are provided in the Appendix (§A).
BM&E: Better Models and Ensembles; HE: Hacking Metrics; RAI: Responsible AI; PM: Predictive Multiplicity; EM: Explanation Multiplicity; OM: Other Multiplicity; RSE: Rashomon Set Exploration; REML: Rashomon Effect in ML; App: Application; Sur: Survey.
2. The Rashomon Effect in Machine Learning
@@
Taking its name from Akira Kurosawa’s 1950 film Rashomon, the Rashomon effect is an epistemological framework that highlights the subjectivity and ambiguity inherent in human perception (Anderson, 2016; Davis et al., 2015; Alia, 1996; Heider, 1988). Borrowing from Davis et al. (2015), the Rashomon effect can be defined as “a combination of a difference of perspective and equally plausible accounts, with the absence of evidence to elevate one above others, […]”. The Rashomon effect has been studied in several different domains, like the influence of cognitive biases on memory (Tindale, 2016; Trabasso, 2018), the impact of culture and the fluidity of truth in ethnographic studies (Heider, 1988), the study of context, medium, and framing of communication (Anderson, 2016; Soreanu and German, 2022), the unreliability of eyewitnesses (Pansky et al., 2005; Hirst and Brown, 2011), and–central to our discussion–algorithmic modelling and machine learning (Breiman, 2001; Black et al., 2022).
@@
The term Rashomon effect was first introduced into algorithmic modelling by Breiman (2001), pointing out the presence of a set of good models that all achieve similar error rates. It has since been used in discussions of statistical modelling (Wang and Tao, 2008; Bonate, 2006; Broadhurst and Kell, 2006; Ueki and Kawasaki, 2013), null hacking (Protzko, 2018), designing robust algorithms (Tulabandhula and Rudin, 2014; Castillo et al., 2008), measuring variable importance (Dong and Rudin, 2019; Fisher et al., 2019), and applications in various domains (Kang et al., 2018; Chantre et al., 2018; Sreedharan et al., 2018). More recently, it has found a resurgence with increasing attention given to multiplicity in machine learning, evident both in studies that directly address the topic (Marx et al., 2020; Black et al., 2022; D’Amour et al., 2022; Rudin et al., 2024; Paes et al., 2023; Del Giudice, 2024; Biecek et al., 2024) and in research that situates multiplicity within the broader context of other fields (Møllersen and Holsbø, 2023; Rudin et al., 2022; Molnar et al., 2020; Jiang et al., 2024b; Biecek and Samek, 2024).
@@
Why do we see the Rashomon effect? In machine learning, data serves as a proxy for the real world, yet it inherently loses information at multiple stages. The first step—translating the world into a data generation process—simplifies complex relationships, introducing randomness to account for uncontrollable aspects. Zhang et al. (2020) termed this “distributional complexity” (associated with ‘aleatoric’–Latin aleatorius–meaning “dice” or “game of chance”), reflecting the challenge of how well the distribution represents the real world. However, even this distribution remains out of our reach; instead, we work with finite samples. This second step of information loss, described as ”approximation complexity” (associated with ‘epistemic’–Latin episteme–meaning “knowledge”) (Zhang et al., 2020), relates to how well can we approximate the underlying generation using finite data. Unlike distributional complexity, which is irreducible, approximation complexity can be mitigated through better data quality and improved algorithms.
@@
Together, distributional and approximation complexities define the fundamental loss of information in machine learning, resulting in gaps where multiple interpretations—and thus the Rashomon effect—can arise. While viewing the Rashomon effect through this lens of information loss is insightful, it does not provide an operational framework for developers to examine or address these challenges. Therefore, in this section, we instead focus on the role of developer choices in model design, laying the foundation for discussions on multiplicity.
Figure 2. Impact of developer choices on individuals and the population downstream.
2.1. Rashomon Effect through Design Choices and Model Selection
@@
Designing a machine learning model involves a series of interconnected or choices. Beginning with the data, decisions are made regarding how to process and filter data, which features to select, etc. (Meyer et al., 2023; Simson et al., 2024; Cavus and Biecek, 2024b) Even the random partitioning of the data into training and validation impacts final model behavior (Friedler et al., 2019). Beyond data, the learning algorithm design further entails numerous decisions: model architecture (Arnold et al., 2024; Rudin et al., 2024), hyperparameters (Bouthillier et al., 2021; Arnold et al., 2024), various forms of stochasticity (Picard, 2021; Bouthillier et al., 2021; Pecher et al., 2024; Sellam et al., [n. d.]; McCoy et al., 2020), and even the evaluation and model selection criteria (Ganesh, 2024; Ganesh et al., 2024). Each decision contributes to the cascade of choices that directly impacts the multiplicity of the trained models (see Figure 2).
@@
Notably, these choices are not always well-informed. In some cases, they are intentional, guided by insights from the literature on the effects of algorithm design on model behaviour (Wu et al., 2021; Ponomareva et al., 2023; Kulynych et al., 2023; Ganesh, 2024). In others, they are conventional, driven by popular trends or convenience (Shwartz-Ziv and Armon, 2022; Dubey et al., 2022; Creel and Hellman, 2022; Sellam et al., [n. d.]). Finally, some choices remain arbitrary, like choosing a random seed. It is through training multiple models and evaluating them, that we grasp the impact of these arbitrary choices. To connect these choices with key subdomains in the multiplicity literature, we introduce the Intent-Convention-Arbitrariness (ICA) framework. More specifically, we argue that these choices exhibit the following properties:
@@
  • Intentional Choices and Steering Model Behaviour: We define certain choices as intentional when developers make deliberate choices with an understanding of their impact on the final model to achieve desired outcomes. Examples include incorporating bias-mitigating regularization to enhance fairness (Hort et al., 2024; Chen et al., 2023), using simpler models for better interpretability (Rudin et al., 2024), or applying data augmentation to improve robustness (Rebuffi et al., 2021; Eghbal-zadeh et al., 2024). These choices are made to steer model behaviour in a particular direction, giving the developer control over navigating the Rashomon set without the need to train multiple models, i.e., avoiding model selection. Intentional choices are typically informed by extensive prior research or other advancements that guide the developer. Note that choices like selecting a pre-trained model because it is the only available option do not qualify as an intentional choice. Although the developer is aware of the impact of their choice, the decision is made out of necessity rather than deliberate intent.
@@
  • Conventional Choices and Homogenization: We define conventional choices as choices made without knowledge of their impact, out of convenience, or due to lack of alternatives. Examples include adopting popular models or hyperparameters without evaluating their suitability for the specific application (Shwartz-Ziv and Armon, 2022; Dubey et al., 2022), such as using neural networks where simpler models would suffice (Rudin, 2019; Shwartz-Ziv and Armon, 2022), or applying out-of-the-box fairness, robustness, or explanation techniques without understanding their implications (Gichoya et al., 2023; Rudin, 2019; Adebayo et al., 2018; Lipton, 2018; Tramer et al., 2020; Carlini et al., 2019). By definition, conventional choices follow established norms or trends within the field rather than addressing specific needs. As a result, these choices contribute to “homogenization”, where models trained by different developers exhibit similar behaviour, and can introduce systemic harm (Creel and Hellman, 2022; Wu et al., 2024; Kleinberg and Raghavan, 2021; Jain et al., 2024b; Bommasani et al., 2022). As shown by Bommasani et al. (2022), even shared components, for instance, common choices across developers, can lead to homogenized outcomes across multiple systems (§6.2).
@@
  • Arbitrary Choices and Model Selection: Unlike intentional or conventional choices, arbitrary choices have an indeterminate relationship with the final model, usually evaluated after training the model. Examples include choosing the random seed for weight initialization, data shuffling, dropout, etc., or arbitrary hyperparameter variations such as selecting between a learning rate of 0.008, 0.01, or 0.012. When dealing with arbitrary choices, making appropriate choices depends on model selection post-training. This has been widely recommended in the literature for auditing multiplicity, creating ensembles, or navigating among competing choices (Kulynych et al., 2023; Long et al., 2024; Creel and Hellman, 2022). However, training multiple models can be expensive, especially for large models or complex hypothesis classes (Hsu et al., 2024, [n. d.]; Kissel and Mentch, 2024; Zhong et al., 2024; Mata et al., 2022; Ciaperoni et al., 2024). Even when feasible, model selection risks overfitting, potentially undermining generalization (Ganesh, 2024; Cooper et al., 2021; Coker et al., 2021). These challenges don’t necessarily dismiss the efficacy of training multiple models. Rather, they highlight the complexities of this approach and motivate a deeper investigation into the benefits and pitfalls of model selection (§5).
In practice, few choices, if any, fall entirely into a single category, Instead, every choice made by a developer involves a mix of intentional, conventional, and arbitrary factors. For instance, consider a developer using dropout testing dropout parameter values 0.1, 0.2, and 0.3. The decision to use dropout is likely intentional, informed by a goal to enhance model robustness. However, the choice of dropout over other regularization techniques may be conventional, driven by its widespread use or integration in popular frameworks. Finally, the specific dropout rates chosen are arbitrary, as their impact on the model’s performance is determined through selection post-training. Understanding the balance between these three factors is essential for navigating the challenges posed by the Rashomon effect.
2.2. Impact of Rashomon Effect

### v2

# Systemizing Multiplicity: The Curious Case of Arbitrariness in Machine Learning

@@ Abstract
Algorithmic modeling relies on limited information in data to extrapolate outcomes for unseen scenarios, often embedding an element of arbitrariness in its decisions. A perspective on this arbitrariness that has recently gained interest is multiplicity—the study of arbitrariness across a set of “good models”, i.e., those likely to be deployed in practice. In this work, we systemize the literature on multiplicity by: (a) formalizing the terminology around model design choices and their contribution to arbitrariness, (b) expanding the definition of multiplicity to incorporate underrepresented forms beyond just predictions and explanations, (c) clarifying the distinction between multiplicity and other lenses of arbitrariness, i.e., uncertainty and variance, and (d) distilling the benefits and potential risks of multiplicity into overarching trends, situating it within the broader landscape of responsible AI. We conclude by identifying open research questions and highlighting emerging trends in this young but rapidly growing area of research.
1 Introduction
@@
Machine learning attempts to approximate the complexities of the world, inevitably simplifying aspects of reality and failing to fully capture its nuances (Hooker 2021; Buolamwini and Gebru 2018; Birhane 2022). It is thus inherently susceptible to arbitrariness, as it attempts to extrapolate outcomes based on limited information. Whether due to imperfect data (Buolamwini and Gebru 2018; Geirhos et al. 2020), flawed modeling assumptions (Hooker 2021; Breiman 2001; Jacobs and Wallach 2021), or inherent unknowability (Wang et al. 2024; Dressel and Farid 2018), arbitrariness is an unavoidable byproduct of data-driven learning. Hence, recognizing and understanding this arbitrariness is crucial for developing responsible learning models.
@@
The study of arbitrariness is not new; it has long been a subject of interest in uncertainty literature, with roots going back centuries in statistics and decision theory (Savage 1972; Bayes 1763; Box and Tiao 1973; Gal 2016). Recently, however, a new paradigm called multiplicity has emerged. First articulated by Breiman (2001), multiplicity has gained popularity due to its focus only on the arbitrariness present within a set of “good models”, i.e., models that pass certain selection criteria and thus are likely to be deployed, known as the Rashomon set. Multiplicity takes an intriguing perspective on arbitrariness in model decisions by instead examining arbitrariness in model selection. Through choices made during development, multiplicity offers an operational lens to the issue of arbitrariness and lays the groundwork for practical solutions in real-world applications.
@@
Figure 1: Systematic review of the number of papers over the years and their categorization. Each paper can have multiple tags, marking all categories of contributions made by the paper. Details of the tags are in the Appendix (§A). BM&E: Better Models & Ensembles; HE: Hacking Metrics; RAI: Responsible AI; PM: Predictive Multiplicity; EM: Explanation Multiplicity; OM: Other Multiplicity; RSE: Rashomon Set Exploration; REML: Rashomon Effect in ML; App: Application; Sur: Survey.
Several works in the literature have provided broad overviews of the field of multiplicity. Black, Raghavan, and Barocas (2022) holds a special place in modern multiplicity literature, offering a discussion of “opportunities”, “concerns”, and potential “solutions” of multiplicity. Similarly, Rudin et al. (2024) discusses several benefits of multiplicity, with a focus on identifying simpler and more interpretable models. At this point, it would be remiss not to acknowledge the dissertations of Semenova (2024); Zhong (2024); Cooper (2024); Hsu (2023); Black (2022); Watson-Daniels (2024); Hasan (2022), contributing valuable perspectives on the role of multiplicity in machine learning. Despite these contributions, the field lacks a systematic review of the literature–clearly needed given its rapid growth (Figure 1). To address this gap, we present the first systematic literature review of multiplicity in machine learning, consolidating existing discussions and identifying overarching trends.
To ensure comprehensive coverage, we search across various online repositories (DBLP & ACM Digital Library) using multiple search terms (‘rashomon’, ‘model multiplicity’, ‘set of good models’), followed by rigorous manual filtering. We eventually found 80 papers that deeply engaged with multiplicity as a central theme in their contributions. Each paper was also manually tagged with all applicable tags, and relevant statistics are presented in Figure 1. The growing interest in the field is evident, with literature on a wide range of problems related to multiplicity. Precise details about the literature review process are in the Appendix (§A).
Contributions. Building on the insights from our literature review, we make the following contributions to multiplicity literature. First, we revisit the Rashomon effect, emphasizing the role of developer choices, and propose a novel Intent-Convention-Arbitrariness (ICA) framework to provide formal foundations for future ethnographic studies (§2). Expanding on this, we extend the definitions of Rashomon sets and multiplicity to include underrepresented and unexplored directions of research, anticipating several future subdomains of multiplicity (§3). Next, we formally distinguish multiplicity from related concepts of uncertainty and variance; and provide both mathematically grounded differences as well as practical guidance on when to adopt each perspective (§4). Finally, we trace two overarching trends in the multiplicity literature: its role in exploring diverse interpretations during model selection (§5), and its broader implications within responsible AI (§6). We conclude by identifying open research questions to encourage future work.
2 The Rashomon Effect in Machine Learning
@@
Taking its name from Akira Kurosawa’s 1950 film Rashomon, the Rashomon effect is an epistemological framework that highlights the subjectivity and ambiguity inherent in human perception (Anderson 2016; Davis, Anderson, and Walls 2015). Borrowing from Davis, Anderson, and Walls (2015), the Rashomon effect can be defined as “a combination of a difference of perspective and equally plausible accounts, with the absence of evidence to elevate one above others, […]”. The Rashomon effect has been studied in several different domains, like the influence of cognitive biases on memory (Tindale 2016; Trabasso 2018), the impact of culture and the fluidity of truth in ethnographic studies (Heider 1988), the study of context, medium, and framing of communication (Anderson 2016; Soreanu and German 2022), the unreliability of eyewitnesses (Pansky, Koriat, and Goldsmith 2005; Hirst and Brown 2011), and–central to our discussion–algorithmic modelling and machine learning (Breiman 2001; Black, Raghavan, and Barocas 2022).
Figure 2: Impact of developer choices on individuals and the population downstream.
@@
The term Rashomon effect was first introduced into algorithmic modelling by Breiman (2001), pointing out the presence of a set of good models that all achieve similar error rates. It has since been used in discussions of statistical modelling (Bonate 2006; Ueki and Kawasaki 2013), null hacking (Protzko 2018), designing robust algorithms (Tulabandhula and Rudin 2014; Castillo et al. 2008), measuring variable importance (Dong and Rudin 2019; Fisher, Rudin, and Dominici 2019), and applications in various domains (Kang et al. 2018; Chantre et al. 2018). More recently, it has found a resurgence with increasing attention given to multiplicity in machine learning, evident both in studies that directly address the topic (Marx, Calmon, and Ustun 2020; Black, Raghavan, and Barocas 2022; Rudin et al. 2024; Del Giudice 2024; Biecek et al. 2024) and in research that situates multiplicity within the broader context of other fields (Møllersen and Holsbø 2023; Rudin et al. 2022; Molnar et al. 2020; Jiang et al. 2024b; Biecek and Samek 2024).
@@
Why do we see the Rashomon effect? In machine learning, data serves as a proxy for the real world, yet it inherently loses information at multiple stages. The first step—translating the world into a data generation process—simplifies complex relationships, introducing randomness to account for uncontrollable aspects. Zhang et al. (2020) termed this “distributional complexity” (associated with ‘aleatoric’–Latin aleatorius–meaning “dice” or “game of chance”), reflecting the challenge of how well the distribution represents the real world. However, even this distribution remains out of our reach; instead, we work with finite samples. This second step of information loss, described as ”approximation complexity” (associated with ‘epistemic’–Latin episteme–meaning “knowledge”) (Zhang et al. 2020), relates to how well we can approximate the underlying generation using finite data. Unlike distributional complexity, which is irreducible, approximation complexity can be mitigated through better data quality and improved algorithms.
@@
Together, distributional and approximation complexities define the fundamental loss of information in learning, resulting in gaps where multiple interpretations, i.e., the Rashomon effect, can arise. While the lens of information loss is insightful, it is limiting in its ability to provide an operational framework to address these challenges. Therefore, we instead focus on the role of developer choices in model design, laying the foundation for discussions on multiplicity.
2.1 Design Choices and Model Selection
@@
Designing a machine learning model involves a series of interconnected choices. Beginning with the data, decisions are made regarding how to process and filter data, which features to select, etc. (Meyer, Albarghouthi, and D’Antoni 2023; Simson, Pfisterer, and Kern 2024; Cavus and Biecek 2024b) Beyond data, the learning algorithm design further entails numerous decisions: model architecture (Arnold et al. 2024; Rudin et al. 2024), hyperparameters (Bouthillier et al. 2021; Arnold et al. 2024), various forms of stochasticity (Picard 2021; Bouthillier et al. 2021; Pecher, Srba, and Bielikova 2024; Sellam et al. 2022; McCoy, Min, and Linzen 2020), and even the evaluation and model selection criteria (Ganesh et al. 2024). Each decision contributes to the cascade of choices that directly impacts the multiplicity of the trained models (see Figure 2).
@@
Notably, these choices are not always well-informed. In some cases, they are intentional, guided by insights from the literature on the effects of algorithm design on model behaviour (Wu et al. 2021; Ponomareva et al. 2023; Ganesh 2024). In others, they are conventional, driven by popular trends or convenience (Shwartz-Ziv and Armon 2022; Dubey, Singh, and Chaudhuri 2022; Creel and Hellman 2022). Finally, some choices remain arbitrary, like choosing a random seed. It is through training multiple models and evaluating them that we grasp the impact of these arbitrary choices. To connect these choices with key subdomains in multiplicity, we introduce the Intent-Convention-Arbitrariness (ICA) framework. More specifically, we argue that these choices exhibit the following properties:
@@
  • Intentional Choices and Steering Model Behaviour. We define intentional choices as deliberate choices made with an understanding of their impact to achieve desired outcomes. Examples include incorporating bias-mitigating regularization to enhance fairness (Hort et al. 2024; Chen et al. 2023), using simpler models for better interpretability (Rudin et al. 2024), or applying data augmentation to improve robustness (Rebuffi et al. 2021; Eghbal-zadeh et al. 2024). These choices steer model behaviour, giving the developer control over navigating the Rashomon set without the need to train multiple models. Intentional choices are typically informed by extensive prior research or other advancements. Note that choices like selecting a pre-trained model because it is the only available option do not qualify as an intentional choice. Although the developer is aware of the impact of their choice, the decision is made out of necessity rather than deliberate intent.
@@
  • Conventional Choices and Homogenization. We define conventional choices as choices made without knowledge of their impact, out of convenience, or due to lack of alternatives. Examples include adopting popular models or hyperparameters without evaluating their suitability for the specific application (Shwartz-Ziv and Armon 2022; Dubey, Singh, and Chaudhuri 2022), such as using neural networks where simpler models would suffice (Rudin 2019; Shwartz-Ziv and Armon 2022), or applying out-of-the-box fairness, robustness, or explanation techniques without understanding their implications (Gichoya et al. 2023; Rudin 2019; Adebayo et al. 2018; Lipton 2018; Tramer et al. 2020; Carlini et al. 2019). By definition, conventional choices follow established norms or trends within the field rather than addressing specific needs. As a result, these choices contribute to “homogenization”, where models trained by different developers exhibit similar behaviour, and can introduce systemic harm (Creel and Hellman 2022; Kleinberg and Raghavan 2021; Bommasani et al. 2022). As shown by Bommasani et al. (2022), even shared components, for instance, some common choices across developers, can lead to homogenized outcomes across multiple systems (§6.2).
    Figure 3: The intent-convention-arbitrariness (ICA) framework with some examples.
@@
  • Arbitrary Choices and Model Selection. Unlike intentional or conventional choices, arbitrary choices have an indeterminate relationship with the final model, usually evaluated after training. Examples include choosing the random seed or arbitrary hyperparameter variations. When dealing with arbitrary choices, making appropriate choices depends on model selection post-training. This has been widely recommended for auditing multiplicity, creating ensembles, or navigating among competing choices (Kulynych et al. 2023; Long et al. 2024; Creel and Hellman 2022). However, training multiple models can be expensive, especially for large models or complex hypothesis classes (Hsu et al. 2024b, a; Kissel and Mentch 2024; Zhong et al. 2024). Even when feasible, model selection risks overfitting, potentially undermining generalization (Ganesh 2024; Cooper et al. 2021). These challenges don’t necessarily dismiss the efficacy of training multiple models. Rather, they highlight the complexities of this approach and motivate a deeper investigation into the benefits and pitfalls of model selection (§5).
In practice, few choices, if any, fall entirely into a single category. Instead, every choice made by a developer involves a mix of intentional, conventional, and arbitrary factors. Understanding this balance is essential for navigating the challenges posed by the Rashomon effect (see Figure 3 for some examples). Our ICA framework also lays a foundation for future ethnographic studies exploring design choices in ML. Understanding the model development culture is essential to uncover homogenization or arbitrariness introduced into the system. Our framework provides both the language and conceptual groundwork necessary for such studies.
2.2 Impact of Rashomon Effect